John is 6 feet tall. At 4 p.m., he cast a shadow that is 8 feet long. At the same time, a nearby light pole cast a shadow that 24 feet long. What is the height of th poll?
Answer by stanbon(75887)

John is 6 feet tall. At 4 p.m., he cast a shadow that is 8 feet long. At the same time, a nearby light pole cast a shadow that 24 feet long. What is the height of the pole?
-----
Use a proportion:
pole/shadow = pole/shadow
---
x/24 = 6/8
-------
x = 24(3/4)
---
x = 18 ft (height of pole
------
Cheers,
Stan H.